# iPrOp: Interactive Prompt Optimization for Large Language Models with a Human in the Loop

### v1

@@ Abstract
Prompt engineering has made significant contributions to the era of large language models, yet its effectiveness depends on the skills of a prompt author. Automatic prompt optimization can support the prompt development process, but requires annotated data. This paper introduces iPrOp, a novel Interactive Prompt Optimization system, to bridge manual prompt engineering and automatic prompt optimization. With human intervention in the optimization loop, iPrOp offers users the flexibility to assess evolving prompts. We present users with prompt variations, selected instances, large language model predictions accompanied by corresponding explanations, and performance metrics derived from a subset of the training data. This approach empowers users to choose and further refine the provided prompts based on their individual preferences and needs. This system not only assists non-technical domain experts in generating optimal prompts tailored to their specific tasks or domains, but also enables to study the intrinsic parameters that influence the performance of prompt optimization. Our evaluation shows that our system has the capability to generate improved prompts, leading to enhanced task performance.
@@ 1 Introduction
However, prompt-based natural language processing (NLP) has demonstrated limited robustness across domains, instances, or label schemes (Plaza-del Arco et al., 2022; Yin et al., 2019; Zhou et al., 2022). It is also challenging to develop reliable methods for evaluation of LLMs that factor in prompt brittleness (Ceron et al., 2024). The question of how to design a well-crafted prompt has received an increasing amount of attention. Although there exists research on analyzing which prompts are more effective for tasks like classification and question answering (Liu et al., 2022; Lu et al., 2022; Xu et al., 2022), the need to efficiently identify high-quality prompts has sparked increased attention into automatic prompt optimization (Shin et al., 2020; Pryzant et al., 2023). However, there is a lack of studies that combines user-guided prompt optimization with data-driven prompt optimization. Given that the user consistites the ultimate authority to develop prompts that satisfy the varying trade-offs across different aspects of a specific task, we consider this an important research gap.
@@
Combining prompt optimization with a user in the loop comes with the potential for a more guided engineering process, from which any user may benefit. Two examples are particularly prominent: (1) Laypeople may require help with prompt development for dedicated tasks. (2) Manual prompt engineering may lead to biased configurations. Prior research has demonstrated the role of human-in-the-loop methodologies in building robust systems across a variety of tasks, including debugging text classifiers (Lertvittayakumjorn et al., 2020), hate speech classification (Kotarcic et al., 2022), and question answering chatbot (Afzal et al., 2024).
@@
To achieve the goal of a system that supports users in their prompt development process, we hypothesize that a set of prompt properties is important to decide if a prompt $p$ is considered better than another prompt $p^{\prime}$. These are (a) the performance of a prompt on some annotated data, for instance measured by F1 (we focus in this paper on text classification tasks); (b) The readability and interpretability of the prompt; (c) The quality of an explanation of the predictions of the prompt; and (d), the alignment of the annotations with the users expectations. We therefore propose an interactive prompt optimization approach with a human-in-the-loop that considers all these aspects. The proposed platform enables studies on the interaction between these various parameters in the spirit of an iterative optimization in which the automatic evaluation of an objective function is supported by a human. We further envision that some decisions may be made automatically, while others require the human to decide on the prompt quality. Such collaborative decision process helps to maintain the high quality of the prompts, while limiting the required user interactions to those of particularly high value.
@@
The repository of the system and an explaination video is available at https://www.uni-bamberg.de/nlproc/ressourcen/iprop/. A screenshot of the user interface is shown in Figure 1.
@@
Figure 2: The conceptual workflow of our iPrOp system. The general workflow is shown in the middle. The left part shows potential human interaction in the various modules. To limit the amount of user interactions, each module can be supported by a simulated interaction.
@@ 2.2 Cooperative Artificial Intelligence
This work is related to the field of cooperative artificial intelligence, which touches upon topics of human-machine interaction and efficient protocols of information exchange, enabling humans to solve tasks collaboratively with machines. Such methods also influenced NLP tasks, such as question answering (Benamara and Saint Dizier, 2003), information retrieval (Manning et al., 2008), and chatbot interactions (Hancock et al., 2019). More recent papers draw their attention on collaborative annotation processes and model direct manipulation (Baur et al., 2020; Wang et al., 2021). However, we introduce “human in the loop” via replacing the automatic evaluation of an objective function by a human.
@@ 2.3 Explainable Artificial Intelligence
Although these tools offer intuitive explanations for classifiers, their reliance on perturbations makes them computationally expensive to apply to LLMs because of the high-dimensional nature and complexity of LLMs. An alternative is to leverage the inherent explainability of LLMs (Mavrepis et al., 2024). Wu et al.’s (2024) analyse of strategies to enhance the transparency of LLMs. Bills et al. (2023) demonstrate that LLMs are able to explain individual neurons in LLMs. This work motivate our attempt to prompt LLMs for the explanations of their predictions.
@@ 3 Methods
Figure 2 visualizes the conceptual workflow of our iPrOp system. The workflow begins with an initial seed prompt and proceeds through iterations of prompt updates and evaluations, led by informative samples, explanations, and data evaluation with performance metrics. To reduce human workload, each step can, in principle, be performed by either the user or the system. To simplify the implementation of our current web interface, the user provides the initial prompt and performs prompt selection, while the system handles the remaining steps.
@@
Here, $M(P)$ is a prompt paraphrasing model that varies the prompts $P$ from the previous iteration. $I(p_{i})$ is a presentation of prompt properties to the user, which consists of
@@
This procedure is also visualized in Figure 2. The initialization of seed prompts ((1) in Figure 2) requires users to describe the task. In simulation scenarios, this process can be substituted with an ontological task description or prompts generated automatically by LLMs. Subsequently, the initial prompts are passed to the optimization modules. In the prompt update module (2a), prompts are paraphrased. As an example, this paraphrasing of ‘Classification task with labels: joy and sadness.’ with a meta-prompt of an LLM ’Rephrase the following prompt’ may lead to ‘Classify the emotion of text into joy and sadness.’. [^1] Our current web interface does not yet support manual prompt adjustments. In the prompt evaluation stage (2b), the human in the loop assesses the prompt quality, as described above. Figure 3 further provides a prototypical display of the relevant information for two prompts to be chosen from. The optimization process is terminated once the user is satisfied (3).
@@
4 System Design and Implementation
Figure 4: The architecture of our iPrOp system composed of the frontend module and the backend module.
The demonstration system includes a web interface for single-label text classification. The architecture of the web interface, as illustrated in Figure 4, is divided into the frontend and the backend modules. The prompt undergoes optimization through multiple iterations between the two modules.
@@ 4 System Design and Implementation
4.1 Frontend
The frontend is implemented using Taipy [^2] https://docs.taipy.io/en/latest/, with a customized interface extending from their LLM Chatbot Demo [^3] https://github.com/Avaiga/demo-chatbot. It receives as user inputs an uploaded dataset, the selection of LLMs, and text describing the task, and displays relevant information about available prompts as outputs.
LLM Selection.
The iPrOp interface allows users to select an LLM, including those accessible via remote API services, such as GPT-4o (OpenAI, 2023) and Claude (Anthropic, 2024), as well as local models, such as LLama3 (Dubey et al., 2024) and Mistral (Jiang et al., 2023). Support for local models is provided through the Ollama server [^4] https://github.com/ollama/ollama and the LangChain Community [^5] https://github.com/langchain-ai/langchain.
Dataset Upload.
To accommodate users’ needs for searching optimal prompts across various domains, the web interface requires users to upload their datasets. The dataset must include a minimum of two columns: one for text and another for the corresponding annotated label. Accepted file formats are .csv, .json, and .jsonl.
User Interaction.
The user enters and sends a texutal task description to the system as an initial prompt. This description must include the names of all possible labels for the task, ensuring consistency with the label values in the uploaded dataset. To select the preferred prompt, they enter the presented prompt id.
System Response.
The user bases the information about the prompt, its explanation on a subset of the training data, the performance score, and example labels. This information is shown as a textual answer in a dialogue manner. The system shows two available prompts with the additional information.
4.2 Backend
The backend generates predictions by prompting an LLM, integrating pre- and post-processing steps prior to and following the LLM interaction.
Pre-processing.
The user-uploaded dataset is divided into training, validation, and testing splits using stratified sampling. Instances shown to users and a subset of data to calculate performance metrics are randomly selected.
Tapping into LLM.
The chosen LLM is used in multiple steps during each iteration. First, the LLM is tasked with adjusting or improving prompts, which are subsequently incorporated into the next iteration. Second, it generates label predictions for text samples. Third, explanations on predicted labels of instances are provided by the LLM.
Post-processing.
Despite specifying in the LLM requests that only the necessary information is to be provided, redundant text and slight format variations may exist. We postprocess the output in a rule-based manner to filter out extraneous content in the LLM output for rephrased prompts and predicted labels. Particularly, we identify true labels in the processed outputs.
@@ 5 Evaluation
We envision our system to enable future research on the interaction of the various aspects to consider when humans make preference decisions on particular prompts under the available information. To validate the principled feasibility of our iPrOp system, we run experiments on three emotion classification datasets using the llama3.1:8b-instruct-fp16 model [^6] https://ollama.com/library/llama3.1:8b-instruct-fp16 (Dubey et al., 2024). In this experiment, we only consider automated classification performance scores and leave an automated evaluation of the other measures or a user study for future work. In this simulation, the prompt is selected corresponding to the weighted F1 score over a fixed subset of the training data. We expect to demonstrate a rising trend during the optimization process to verify the effectiveness of our system.
@@
We select three datasets for single labeled emotion classification task from Bostan and Klinger (2018), namely TEC, which covers general topics on tweets (Mohammad, 2012); Grounded-Emotions, which focuses on event-related topics on tweets; and Tales-Emotion, which is built upon fairytales (Alm and Sproat, 2005).
@@
Figure 5, illustrates the F1 scores over 15 iterations. We observe an overall increasing trend in both training and validation data.
@@ 6 Conclusions and Future Work
In this paper, we proposed an interactive prompt optimization system that incorporates human intervention for finding well-performing prompts, based on various information that is destilled from the prompt and its performance on some user-provided data. With this system, we suggested to aggregate information that may be relevant for users to decide on prompt preferences. In our evaluation, we focused on the performance of the predictions alone, but the important aspect of this work is that it enables future research on the interaction of explanations, prompt performance, example predictions, and the prompt readibility. These studies need to be performed both from an automated angle as well as in user studies.
The proposed system has revealed several challenges that deserve further investigation. There is a need to explore more effective methodologies for enhancing the diversity of rephrased prompts. It is important to limit the numbers of instances shown to the user, and that selection requires methods to do so. It is essential to optimize the various meta-prompts in the system. Additionally, the optimization algorithm is essential to improving the efficiency and user-friendliness of our system.
@@ Acknowledgments
This paper is supported by the project INPROMPT (Interactive Prompt Optimization with the Human in the Loop for Natural Language Understanding Model Development and Intervention, funded by the German Research Foundation, KL 2869/13-1, project numer 521755488).
Limitations
Although the iPrOp system offers a convenient web interface for non-technical users to attain suitable prompts, it has several limitations that warrant consideration in the future enhancement. First, because of the system’s inherent design as a guidance tool for prompt optimization, it does not exhibit the open interactivity of a chatbot. Instead, users are expected to adhere closely to the system’s instructions throughout the optimization process. Second, as a primary web interface for presenting our system, we do not provide any tools of user account or data management, while only simple data storage is integrated. Third, developing an effective strategy to address problems related to train-validation-test splitting for user-provided datasets of varying sizes remains an ongoing challenge. Lastly, the development of prompt optimization iterations partially depends on the quality and variability of prompt rephrasing. This implies that rephrased prompts may occasionally retain low quality across multiple iterations. Furthermore, we observe that certain datasets exhibit limited sensitivity to divergent prompts, allowing a simple or even naive initial prompt to achieve superior performance.
@@
Our system is designed with careful attention to ethical standards in data usage, privacy, and compliance with the ACL Code of Ethics. Our method does not contribute to the republication or redistribution of any datasets. The datasets used for testing and evaluation are publicly available and we ensure that they have been collected according to ethical standards before using them. To safeguard user privacy, all data provided by users is stored exclusively on their local machines. While potential risks associated with the underlying LLMs could result in the exposure of user-provided datasets, we aim to mitigate these risks by offering more secure local models. In addition, our system cannot guarantee that the optimal prompts identified are state of the art for specific tasks. Furthermore, individual preferences may introduce biases, which could potentially mislead users. We are committed to continuously monitoring and improving the ethical performance of our system.

### v2

@@ Abstract
Prompt engineering has made significant contributions to the era of large language models, yet its effectiveness depends on the skills of a prompt author. This paper introduces iPrOp, a novel interactive prompt optimization approach, to bridge manual prompt engineering and automatic prompt optimization while offering users the flexibility to assess evolving prompts. We aim to provide users with task-specific guidance to enhance human engagement in the optimization process, which is structured through prompt variations, informative instances, predictions generated by large language models along with their corresponding explanations, and relevant performance metrics. This approach empowers users to choose and further refine the prompts based on their individual preferences and needs. It can not only assist non-technical domain experts in generating optimal prompts tailored to their specific tasks or domains, but also enable to study the intrinsic parameters that influence the performance of prompt optimization. The evaluation shows that our approach has the capability to generate improved prompts, leading to enhanced task performance.
@@ 1 Introduction
However, prompt-based natural language processing (NLP) has demonstrated limited robustness across domains, instances, or label schemes (Plaza-del Arco et al., 2022; Yin et al., 2019; Zhou et al., 2022). It is also challenging to develop reliable methods for evaluation of LLMs that factor in prompt brittleness (Ceron et al., 2024). The question of how to design a well-crafted prompt has received an increasing amount of attention. Although there exists research on analyzing which prompts are more effective for tasks like classification and question answering (Liu et al., 2022; Lu et al., 2022; Xu et al., 2022), the need to efficiently identify high-quality prompts has sparked increased attention into automatic prompt optimization (Shin et al., 2020; Pryzant et al., 2023). However, they tend to overlook the inherent contextuality and the domain-dependent nature of prompt engineering (Pei et al., 2025; Anthropic, 2024). There is a lack of studies that combines user-guided prompt optimization with data-driven prompt optimization. Given that the user constitutes the ultimate authority to develop prompts that satisfy the varying trade-offs across different aspects of a specific task, we consider this an important research gap.
@@
Combining prompt optimization with a user in the loop comes with the potential for a more guided engineering process, from which any user may benefit. Two examples are particularly prominent: (1) Technical laypeople may require help with prompt development for dedicated tasks. (2) Manual prompt engineering may lead to biased configurations, as generic prompts often fail to capture the complexities and nuances specific to particular domains, such as medical knowledge (Lu et al., 2023). Prior research has demonstrated the role of human-in-the-loop methodologies in building robust systems across a variety of tasks, including debugging text classifiers (Lertvittayakumjorn et al., 2020), hate speech classification (Kotarcic et al., 2022), and question answering chatbots (Afzal et al., 2024).
@@
To achieve the goal of supporting users in their prompt development process, we hypothesize that a set of prompt properties is important to decide if a prompt $p$ is considered better than another prompt $p^{\prime}$. These are (a) the performance of a prompt on some annotated data, for instance measured by F1 (we focus in this paper on text classification tasks); (b) The readability and interpretability of the prompt; (c) The quality of an explanation of the predictions of the prompt; and (d), the alignment of the annotations with the users expectations. We therefore propose an interactive prompt optimization approach with a human-in-the-loop that considers all these aspects. The proposed approach enables studies on the interaction between these various parameters in the spirit of an iterative optimization in which the automatic evaluation of an objective function is supported by a human. We further envision that some decisions may be made automatically, while others require the human to decide on the prompt quality. Such collaborative decision process helps to maintain the high quality of the prompts, while limiting the required user interactions to those of particularly high value.
@@
The repository of a prototypical web interface for the iPrOp approach and an explanation video is available at https://www.uni-bamberg.de/nlproc/ressourcen/iprop/. Figure 1 presents a screenshot of the web-based user interface.
@@
Figure 2: The conceptual workflow of our iPrOp approach. The general workflow is shown in the middle. The left part shows potential human interaction in the various modules. To limit the amount of user interactions, each module can be supported by a simulated interaction.
@@ 2.2 Cooperative Artificial Intelligence
This work is related to the field of cooperative artificial intelligence, which touches upon topics of human-machine interaction and efficient protocols of information exchange, enabling humans to solve tasks collaboratively with machines. Such methods also influenced NLP tasks, such as question answering (Benamara and Saint Dizier, 2003), information retrieval (Manning et al., 2008), and chatbot interactions (Hancock et al., 2019). More recent papers draw their attention on collaborative annotation processes and model direct manipulation (Baur et al., 2020; Wang et al., 2021). However, we introduce a human-in-the-loop via replacing the automatic evaluation of an objective function by a human. Prior research has explored incorporated human feedback by presenting users with responses generated from paired prompts and asking for their preferences (Lin et al., 2024). In contrast, our framework offers a more comprehensive structure, encompassing a broader range of factors that should be considered during human evaluation.
@@ 2.3 Explainable Artificial Intelligence
Although these tools offer intuitive explanations for classifiers, their reliance on perturbations makes them computationally expensive to apply to LLMs because of the high-dimensional nature and complexity of LLMs. An alternative is to leverage the inherent explainability of LLMs (Mavrepis et al., 2024). Wu et al.’s (2024) analysis of strategies to enhance the transparency of LLMs. Bills et al. (2023) demonstrate that LLMs are able to explain individual neurons in LLMs. This work motivates our attempt to prompt LLMs for the explanations of their predictions.
@@ 3 Methods
Figure 2 visualizes the conceptual workflow of our iPrOp approach. The workflow begins with an initial seed prompt and proceeds through iterations of prompt updates and evaluations, led by informative samples, explanations, and data evaluation with performance metrics. To reduce human workload, each step can, in principle, be performed either by the user or automatically.
@@
Here, $M(P)$ is a prompt paraphrasing model that varies the prompts $P$ selected from the previous iteration. $I(p_{i})$ is a presentation of prompt properties to the user, which consists of
@@
This procedure is also visualized in Figure 2. The initialization of seed prompts ((1) in Figure 2) requires users to describe the task. In simulation scenarios, this process can be substituted with an ontological task description or prompts generated automatically by LLMs. Subsequently, the initial prompts are passed to the optimization modules. In the prompt update module (2a), prompts are paraphrased. As an example, this paraphrasing of ‘Classification task with labels: joy and sadness.’ with a meta-prompt of an LLM ‘Rephrase the following prompt’ may lead to ‘Classify the emotion of text into joy and sadness.’
In the prompt evaluation stage (2b), the human in the loop assesses the prompt quality, as described above. Figure 3 further provides a prototypical display of the relevant information for two prompts to be chosen from. In the current prototype interface, the explanations are automatically generated by prompting a LLM. For instance, the specific prompt used is: ‘In your answer, provide only the label you choose and the explanation of your choice.’. Examples of the generated explanations during the evaluation process are provided in the Appendix A. The optimization process is terminated once the user is satisfied (3).
@@ 4 Evaluation
We envision our iPrOp approach to enable future research on the interaction of the various aspects to consider when humans make preference decisions on particular prompts under the available information. To validate the principled feasibility of our approach, we run experiments on three emotion classification datasets using the llama3.1:8b-instruct-fp16 model [^1] https://ollama.com/library/llama3.1:8b-instruct-fp16 (Dubey et al., 2024). In this experiment, we only consider automated classification performance scores and leave an automated evaluation of the other measures or a user study for future work. In this simulation, the prompt is selected corresponding to the weighted F1 score over a fixed subset of the training data. We expect to demonstrate a rising trend during the optimization process to verify the effectiveness of our approach.
@@
We select three datasets for single labeled emotion classification task from Bostan and Klinger (2018), namely TEC, covering general topics on tweets (Mohammad, 2012); Grounded-Emotions, focusing on event-related topics on tweets (Liu et al., 2017); and Tales-Emotion, built upon fairytales (Alm and Sproat, 2005).
@@
Figure 4 illustrates the F1 scores over 15 iterations. We observe an overall increasing trend in both training and validation data.
@@ 5 Conclusions and Future Work
We proposed interactive prompt optimization as a novel approach to configure instruction-tuned language models. The user is guided by information that is distilled from the prompt and its performance on user-provided data. With this approach, we suggested to aggregate information that may be relevant for users to decide on prompt preferences.
The proposed approach has revealed several challenges that deserve further investigation. There is a need to explore more effective methodologies for enhancing the diversity of rephrased prompts. It is important to limit the numbers of instances shown to the user, and that selection requires methods to do so. It is essential to optimize the various meta-prompts in the approach. Additionally, the optimization algorithm is essential to improving the efficiency and user-friendliness of our approach.
We envision that our iPrOp approach lays the groundwork for future research by addressing several open questions: (Q1) Which parameters do influence the performance of the workflow configuration in this approach? We presume that the example selection to better understand how the prompt performs affects a user’s ability to estimate which prompt is preferable. Further, the methods to explain the prompt prediction are crucial. Finally, underlying aspects such as the model and its robustness are relevant factors for the approach to succeed. (Q2) How do prompts evolve throughout the optimization iterations? An aspect of this question is what is the difference between automatic prompt optimization and the human optimization is, and in which cases the human intervention is indeed helpful. (Q3) To what extent can human involvement be reduced while maintaining a balanced trade-off across competing evaluation criteria? Can the interactive prompt optimization approach be a collaborative learning procedure, in which the machine only requests information if needed? We propose to study these research questions based on the paradigm of interactive prompt optimization introduced in this paper.
Limitations
Although the iPrOp approach offers a convenient interface for non-technical users to attain suitable prompts, it has several limitations that warrant consideration in the future enhancement. First, in an effort to provide comprehensive explanations of LLM predictions, the challenge of computation time remains significant, and as a result, the streaming output is not effectively communicated to users. Second, developing an effective strategy to address problems related to train-validation-test splitting for user-provided datasets of varying sizes remains an ongoing challenge. Third, the development of prompt optimization iterations partially depends on the quality and variability of prompt rephrasing. This implies that rephrased prompts may occasionally retain low quality across multiple iterations. Furthermore, we observe that certain datasets exhibit limited sensitivity to divergent prompts, allowing a simple or even naive initial prompt to achieve superior performance.
@@ Acknowledgments
This paper is supported by the project INPROMPT (Interactive Prompt Optimization with the Human in the Loop for Natural Language Understanding Model Development and Intervention, funded by the German Research Foundation, KL 2869/13-1, project number 521755488).
@@
Our approach is designed with careful attention to ethical standards in data usage, privacy, and compliance with the ACL Code of Ethics. Our method does not contribute to the republication or redistribution of any datasets. The datasets used for testing and evaluation are publicly available and we ensure that they have been collected according to ethical standards before using them. To safeguard user privacy, all data provided by users is stored exclusively on their local machines. While potential risks associated with the underlying LLMs could result in the exposure of user-provided datasets, we aim to mitigate these risks by offering more secure local models. In addition, our approach cannot guarantee that the optimal prompts identified are state of the art for specific tasks. Furthermore, individual preferences may introduce biases, which could potentially mislead users. We are committed to continuously monitoring and improving the ethical performance of our approach.
Appendix A Appendix
Text Predicted Label & Explanation
hey you! Hope school is going well for you! Love $\heartsuit$ Joy. The presence of the $\heartsuit$ symbol and the affectionate message ("Love ya") convey a positive emotional tone, indicating joy.
Mmmm breakfast in bed :) Joy. The person is expressing excitement and delight through the use of a happy emoticon :).
A wonderful story about how an act of kindness can change someone’s life. Joy. The use of "wonderful" to describe the story suggests a positive tone, and the sharing of it on social media implies that the act of kindness is uplifting and inspiring.
Someday…. Sadness. The use of the word "someday" can imply a sense of longing or wistfulness, and the fact that itś used in conjunction with two links to online content may suggest that the author is reflecting on something theyv́e written in the past.
With immigration drop, who needs a wall? Sadness. The text mentions an "immigration drop", which implies a negative change or decrease in something related to immigration. This is likely to evoke feelings of sadness or concern rather than joy. The use of the phrase "who needs a wall" also suggests a sense of pessimism or hopelessness, further indicating that the overall tone of the text is more negative and therefore classified as sadness.
do not try to stop them Sadness. The use of "do not try to stop them" is a warning or cautionary statement, often implying something bad or undesirable is happening. The link provided does not provide context, but the tone of the message suggests a sense of urgency and possibly desperation, which are common indicators of sadness or concern.
Table 1: Examples of an explanation generated for a predicted label in the Grounded-Emotions dataset task.